A bomb calorimeter is filled with 4.4 atm of an unknown gaseous hydrocarbon fuel and an excess of oxygen gas at a temperature of 21 °C. The bomb calorimeter has a volume of 5 L and a heat capacity of 8.34 kJ/K. After the combustion reaction is complete the final temperature of the calorimeter is found to be 231.6 °C. Answer the following questions based on this information.

(a) What amount of heat in kJ was released from the combustion reaction of the unknown fuel? Please give your answer to 2 decimal places.



(b) What is the molar heat of combustion for the reaction in kJ/mol? Please give your answer to 2 decimal places.


(c) Given the table of data below, what is the identity of the unknown gas?

Fuel Chemical Formula Molar Heat of Combustion
hydrogen H2 -285.9 kJ/mol
propene CH2CHCH3 -1927.3 kJ/mol
ethene CH2CH2 -1324.1 kJ/mol
cyclopropane C3H6 -2078.8 kJ/mol



The identity of the unknown gas is  hydrogen propene ethene cyclopropane ?


(d) What is the heat of combustion for the reaction in kJ/g? Please give your answer to 2 decimal places.
Answer #1

a)

heat capacity = 8.34 kJ/ K

temperature difference = 231.6 - 21 = 210.6 K

heat released = Cp dT

                      = 8.34 x 210.6

                      = 1.76 x 10^3 kJ

b)

P V = n R T

4.4 x 5 = n x 0.0821 x 294

n = 0.911 mol

heat of combustion = Q / n

                               = 1.76 x 10^3 / 0.911

                              = -1.93 x 10^3 kJ/mol

c)

the unknown hydrocarbon = propene

d)

moles = 0.911

mass = 0.911 x 42.02 = 38.33 g

heat of combustion = 1756 / 38.33

                              = - 45.81 kJ/g